The length of a rectangle is 3 feet more than twice it's width. Find the length, perimeter, and area of a rectangle a) if it's width is 5; b) if it's width is 10
Answer by palanisamy(496)

a) Let the width of the rectangle = 5 feet
Then, its length = 5+3 = 8
Perimeter of the rectangle = 2*5+2*8
= 10+16
= 26 feet
Area of the rectangle = 5*8
= 40sq feet
b)Let the width of the rectangle = 10 feet
Then, its length = 10+3 = 13
Perimeter of the rectangle = 2*10+2*13
= 20+26
= 46 feet
Area of the rectangle = 10*13
= 130sq feet